word problem

In the 2007 regular baseball season , the Boston Red Sox won 36 less than twice as many games as they lost . They played 162 regular season games. How many wins and losses did the team have ?

Answer by simranb(63)

Let the lost sessions be x
Then won sessions=2x-36
Total no. of sessions=162
Equation
(x)+(2x-36)=162
x+2x-36=162
3x-36=162
3x=162+36
3x=198
x=198/3
x=66
So,
Losses=66
wins=2x-36
=(2*66)-36
=132-36
=96
Hope this helps!